help me solve this question ytan⁡xdy/dx=(4+y^2 ) sec^2⁡x thanks

Grade:1

1 Answers

Harshit Singh
3 years ago
Dear Student

On separating variables ;
y dy/(4+y2) = (sec2x/tanx)dx
on integrating both sides
let y^2 =m tan x = n
ydy = 0.5*dm sec^2x dx= dn
on integrating
(1/2)log(4+y2) = log (tan x) +c

Thanks